A photon has an energy of 5.53 × 10 -17 J. What is its frequency in s -1 (h = 6.63 x 10 -34 J s)?
1
Expert's answer
2019-09-25T05:46:18-0400

E= hv. v= E/h= 5.53 × 10 -17 / 6.63 x 10 -34= 8.34087481 x 10 -16